# Unanswerable Visual Question Answering

### v1

@@ Abstract
Teaching Visual Question Answering (VQA) models to abstain from unanswerable questions is indispensable for building a trustworthy AI system. Existing studies, though have explored various aspects of VQA, yet marginally ignored this particular attribute. This paper aims to bridge the research gap by contributing a comprehensive dataset, called UNK-VQA. The dataset is specifically designed to address the challenge of questions that can be unanswerable. To this end, we first augment the existing data via deliberate perturbations on either the image or question. In specific, we carefully ensure that the question-image semantics remain close to the original unperturbed distribution. By means of this, the identification of unanswerable questions becomes challenging, setting our dataset apart from others that involve mere image replacement. We then extensively evaluate the zero- and few-shot performance of several emerging multi-modal large models and discover significant limitations of them when applied to our dataset. Additionally, we also propose a straightforward method to tackle these unanswerable questions. This dataset, we believe, will serve as a valuable benchmark for enhancing the abstention capability of VQA models, thereby leading to increased trustworthiness of AI systems. We have made the dataset available to facilitate further exploration in this area.
@@
Visual Question Answering, Unanswerable Questions, Vision and Language
@@ I Introduction
Visual Question Answering (VQA) serves as a fundamental task in the pursuit of achieving artificial general intelligence. Conventional approaches often view VQA as a multi-category classification problem, wherein each class aligns with a frequent answer in the dataset under consideration [1, 2, 3]. These benchmark datasets encapsulate a spectrum of intricate vision-language comprehension abilities, such as out-of-distribution generalization [3], compositional reasoning [4], and utilization of external knowledge [5, 6].
@@
Despite the crucial function of these benchmarks during the early phase of exploration, recent progress has largely been driven by the emergence of large models [7, 8, 9]. These large models overshadow the common training-and-evaluation paradigm in various domains, including VQA. For example, GPT-4 [8] achieves a zero-shot accuracy of $\sim$80on the VQA v2 dataset [1], significantly outperforming most supervised approaches. Pushing more limits on these traditional well-studied benchmarks thus appears to yield marginal contributions. While effective, these large foundation models, however, are notoriously criticized for their inability of trustworthy. One demonstration is that they are less capable of abstaining from answering questions that cannot be answered or are beyond their scope of knowledge [10].
@@
This paper advocates a dataset that teaches machines to identify and reject unanswerable visual questions. Arguably, annotating a large-scale dataset with unanswerable questions is extremely laborious and difficult to control. To address this challenge, we suggest introducing perturbations to existing data and curating hard instances that can potentially deceive models. Based on the most popular VQA v2 dataset [1], we apply five different types of nuanced perturbations to either the given image or the question, as shown in Fig. 1. To assure high quality, we engage the Amazon Mechanical Turk (MTurk) workers to label them manually. In short, the total number of participating workers is $>$4,000, and each instance is annotated by a minimum of three workers. This finally amounts to 10K instances in our dataset, which we name as UNK-VQA. As can be observed from Fig. 1, masking the most relevant object (I-2) often leads to more difficulties in performing question answering. Unlike existing approaches [11, 12, 13] that employ unsupervised learning techniques to address unanswerable visual questions, our proposed dataset enables model training towards more generalizable VQA methods.
@@
We justify the performance of multiple multi-modal large models including Otter [14], Open-Flamingo [9], and InstructBLIP [15]. Our findings indicate that: 1) In terms of zero-shot performance, our UNK-VQA dataset reveals limitations in these large models, unlike the impressive results observed in previous general VQA datasets. 2) For few-shot performance, when we introduce more instances to these models, consistent improvements in their performance can be observed. Additionally, we also propose a simple method to enhance VQA models with the ability of abstention. The key to our method lies in the design of selection functions, which we have developed in several variants, such as a binary classifier and entropy-based methods. Considering the resource-intensive nature of fine-tuning large models, we apply this method to several conventional and recent medium-sized VL Transformer models.
@@
This dataset reflects the fact that existing multi-modal large models, in comparison to their counterparts in the language domain, are not as omnipotent. It sheds light on various issues beyond the challenge of abstaining from answering unanswerable questions. Problems such as consistency in VQA [16], factuality, and hallucination [8] further emphasize the need for extensive research in this domain.
@@
Figure 2: Illustration of two text-based perturbations. We utilize a strong multi-modal model, i.e., BLIP [17] to generate answers (Base-A) to the perturbed question (Pert-Q). In the case of word replacement examples, when we replace the anchor noun word with an alternative, the baseline still generates a reasonable answer, even though the modified question becomes unanswerable. Regarding the semantic negation examples, each question can have uncounted potential answers.
@@ II-A Perturbation Procedure and Quality Control
For a given question $Q$, we first leverage NLTK toolkits [^1] https://www.nltk.org/. to detect nouns and select one noun as the anchor word $w_{c}$. After that, we estimate the proximity between $w_{c}$ and the remaining words in Glove using their pre-trained word embeddings [18]. The $k$ nearest neighboring words, denoted as $\mathcal{W}_{c}$, serve as candidates for replacement. We then enhance the original question by generating augmented questions $\mathcal{Q}_{c}$. This is done by replacing $w_{c}$ with each $w_{i}\in\mathcal{W}_{c}$. One instance is shown in Fig. 2 that the anchor word bridge is replaced with tunnel. We believe that these augmented questions are challenging to answer. During this procedure, we eliminate duplicates through lexical rules such as plural and tense comparisons. Nevertheless, we have observed that some questions with replaced words rarely occur naturally. This potential shortcut may easily lead the perturbed question to be unanswerable based on semantic coherence alone. To approach this, we employ a large language model (LM), i.e., GPT-2-Large [19], to filter out examples with high perplexity after augmentation. The resulting set of questions that pass through the LM filter is defined as:
@@
where $w_{i}$ represents the $i$-th token of $Q$, $\mathbf{\Theta}$ is the parameter of the pre-trained language model, and $\epsilon$ is a pre-defined threshold and we empirically set it to 0.4.
Semantic Negation (T-2) Another commonly used approach in text-based perturbation is the semantic negation of the original question. To ensure the perturbation is effective, we carefully exclude questions that have binary answers to avoid trivial errors. In practice, we initially perform dependency parsing to identify all verbs and auxiliaries in the question. We then add negation to these words, such as converting copulas to their negated form or simply adding did not before the verbs. If none of the above steps are applicable, we check if any negation keywords like not, hardly, or never, have already been mentioned in the question, and remove them accordingly. However, it is important to note that using these heuristic rules for negation can sometimes lead to a significant loss in semantic coherence and fluidity. Similar to the approach used in word replacement, we also rely on a language model to ensure the quality of the synthesized text.
@@
To achieve this, we leverage a powerful pre-trained vision encoder, CLIP [20], to extract image features. These features are then used to rank all available images based on their similarity to the anchor image, resulting in a selection of the top 50 candidates that share similar semantics with the given image $I$. Next, we proceed to detect the objects in all the images, as well as the conceptual information within the questions, especially the answers. Our objective is to identify and remove images that have a high degree of overlap with the concepts mentioned in the questions and answers. To quantify this overlap, we define a semantic overlap score $s_{op}$:
@@ III-A Evaluation Setup
We adopt four multi-modal large models for this evaluation.
@@
  • Otter-MPT-7B [14] shares a similar model architecture to Open-Flamingo but has been specifically optimized on the Multi-Modal In-Context Instruction Tuning (MIMIC-IT) dataset. This optimization aims to enhance the model’s ability to follow the in-context examples accurately for better few-shot inference.
@@
#Shots #Ans #Una Acc(%$\uparrow$ OoS(%$\downarrow$
@@ VI-A Visual Question Answering Datasets
VQA contributes an essential ingredient to the compelling ‘AI-Complete’ tasks. At its early emerging stage, significant efforts were dedicated to constructing a general VQA dataset, aiming to encompass all aspects of visual reasoning from textual inputs [37, 38, 39, 40]. For instance, DAQUAR [41] and COCO-QA [42] employe automated question-answer pair generation techniques, minimizing the need for extensive human labor. Visual Madlibs [43] utilizes fill-in-the-blank templates to generate VQA instances. Additionally, benchmark datasets such as [37, 44] collected large-scale human-annotated VQA data, including both natural images and abstract scenes. With increasing attention to the VQA task, the model performance on these pioneering datasets gradually approached saturation with respect to human performance.
@@ VII Conclusion and Future Work
This paper advocates a novel dataset to enable VQA models with the ability of abstention. By leveraging this dataset, we probe the robustness of multi-modal large models and identify their relatively less satisfactory performance. We then propose a straightforward approach to help train VQA models with the inclusion of unanswerable questions. We evaluate the effectiveness of this approach by integrating it into several baseline methods. The findings reported in this paper highlight two promising future research directions:
@@
2) Exploring additional possibilities for multi-modal large model training. The existing VL large model is still in its infancy, with most efforts focused on aligning them with LLMs. However, we have discovered that these large models exhibit significant limitations when it comes to tasks such as instruction following and visual understanding. As such, a more promising direction could involve building multi-modal large models that leverage the intrinsic attributes of multi-modality, rather than relying solely on mapping it back to language.

### v3

# UNK-VQA: A Dataset and A Probe into Multi-modal Large Models’ Abstention Ability

@@ Abstract
Teaching Visual Question Answering (VQA) models to refrain from answering unanswerable questions is necessary for building a trustworthy AI system. Existing studies, though have explored various aspects of VQA but somewhat ignored this particular attribute. This paper aims to bridge the research gap by contributing a comprehensive dataset, called UNK-VQA. The dataset is specifically designed to address the challenge of questions that models do not know. To this end, we first augment the existing data via deliberate perturbations on either the image or question. In specific, we carefully ensure that the question-image semantics remain close to the original unperturbed distribution. By this means, the identification of unanswerable questions becomes challenging, setting our dataset apart from others that involve mere image replacement. We then extensively evaluate the zero- and few-shot performance of several emerging multi-modal large models and discover their significant limitations when applied to our dataset. Additionally, we also propose a straightforward method to tackle these unanswerable questions. This dataset, we believe, will serve as a valuable benchmark for enhancing the abstention capability of VQA models, thereby leading to increased trustworthiness of AI systems. We have made the dataset available to facilitate further exploration in this area.
@@
Visual Question Answering, Unanswerable Questions, Multi-modal Large Models
@@ I Introduction
Visual Question Answering (VQA) serves as a fundamental task in the pursuit of achieving artificial general intelligence. Conventional approaches often view VQA as a multi-category classification problem, wherein each class aligns with a frequent answer in the dataset under consideration [1, 2, 3]. These benchmark datasets encompass a spectrum of intricate vision-language comprehension abilities, such as out-of-distribution generalization [3], compositional reasoning [4], and utilization of external knowledge [5, 6].
@@
Despite the crucial function of these benchmarks during the early phase of exploration, recent progress has largely been driven by the emergence of large models [7, 8, 9]. These large models overshadow the common training-and-evaluation paradigm in various domains, including VQA. For example, GPT-4 [8] achieves a zero-shot accuracy of $\sim$80on the VQA v2 dataset [1], significantly outperforming most supervised approaches. Pushing the limits on these traditional well-studied benchmarks thus appears to yield marginal contributions. While effective, these large foundation models, however, are notorious for being trustworthy. One demonstration is that they are less capable of abstaining from answering questions that cannot be answered or are beyond their scope of knowledge [10].
@@
This paper proposes a dataset that teaches machines to identify and refuse unanswerable visual questions. Arguably, annotating a large-scale dataset with unanswerable questions is extremely laborious and difficult to control. To address this challenge, we suggest introducing perturbations to existing data and curating hard instances that can potentially deceive models. Based on the most popular VQA v2 dataset [1], we apply five different types of nuanced perturbations to either the given image or the question, as shown in Fig. 1. To assure high quality, we engage the Amazon Mechanical Turk (MTurk) workers to label them manually. In short, the total number of participating workers is $>$4,000, and each instance is annotated by a minimum of three workers. This finally amounts to 10K instances in our dataset, which we name as UNK-VQA. As can be observed from Fig. 1, masking the most relevant object (I-2) often leads to more difficulties in performing question answering. Unlike existing approaches [11, 12, 13] that employ unsupervised learning techniques to address unanswerable visual questions, our proposed dataset enables model training towards more generalizable VQA methods.
@@
We evaluate the performance of multiple multi-modal large models including Otter [14], Open-Flamingo [9], and InstructBLIP [15]. Our findings indicate that: 1) In terms of zero-shot performance, our UNK-VQA dataset reveals limitations in these large models, unlike the impressive results observed in previous general VQA datasets. 2) For few-shot performance, when we introduce more instances to these models, consistent improvements in their performance can be observed. Additionally, we also propose a simple method to enhance VQA models with the ability of abstention. The key to our method lies in the design of selection functions, which we have developed in several variants, such as a binary classifier and entropy-based methods. Considering the resource-intensive nature of fine-tuning large models, we apply this method to several conventional and recent medium-sized VL Transformer models.
@@
This dataset reflects the fact that existing multi-modal large models, in comparison to their counterparts in the language domain, are not as omnipotent. It sheds light on various issues beyond the challenge of abstaining from answering unanswerable questions. Problems such as consistency in VQA [16], factuality, and hallucination [8] further emphasize the need for further research in this domain.
@@
Figure 2: Illustration of two text-based perturbations. We utilize a strong multi-modal model, i.e., BLIP [17] to generate answers (Base-A) to the perturbed question (Pert-Q). In the case of word replacement examples, when we replace the anchor noun word with an alternative, the baseline still generates a reasonable answer, even though the modified question becomes unanswerable. Regarding the semantic negation examples, each question can have an uncoutable number of potential answers.
@@ II-A Perturbation Procedure and Quality Control
For a given question $Q$, we first employ the NLTK toolkit [^1] https://www.nltk.org/. to detect nouns and select one noun as the anchor word $w_{c}$. After that, we estimate the proximity between $w_{c}$ and the remaining words in Glove using their pre-trained word embeddings [18]. The $k$ nearest neighboring words, denoted as $\mathcal{W}_{c}$, serve as candidates for replacement. We then enhance the original question by generating augmented questions $\mathcal{Q}_{c}$. This is done by replacing $w_{c}$ with each $w_{i}\in\mathcal{W}_{c}$. One instance is shown in Fig. 2 that the anchor word bridge is replaced with tunnel. We believe that these augmented questions are challenging to answer. During this procedure, we eliminate duplicates through lexical rules such as plural and tense comparisons. Nevertheless, we have observed that some questions with replaced words rarely occur naturally. This potential shortcut may easily lead the perturbed question to be unanswerable based on semantic coherence alone. To approach this, we employ a large language model (LM), i.e., GPT-2-Large [19], to filter out examples with high perplexity after augmentation. The resulting set of questions that pass through the LM filter is defined as:
@@
where $w_{i}$ represents the $i$-th token of $Q$, $\mathbf{\Theta}$ is the parameter set of the pre-trained language model, and $\epsilon$ is a pre-defined threshold and we empirically set it to 0.4.
Semantic Negation (T-2) Another commonly used approach in text-based perturbation is the semantic negation of the original question. To ensure the perturbation is effective, we carefully exclude questions that have binary answers to avoid trivial errors. In practice, we initially perform dependency parsing to identify all verbs and auxiliaries in the question. We then add negation to these words, such as converting copulas to their negated form or simply adding did not before the verbs. If none of the above steps are applicable, we check if any negation keywords like not, hardly, or never, have already been mentioned in the question, and remove them accordingly. However, it is important to note that using these heuristic rules for negation can sometimes lead to a significant loss in semantic coherence and fluidity. Similar to the approach used in word replacement perturbation, we also rely on a language model to ensure the quality of the synthesized text.
@@
To achieve this, we utilize a powerful pre-trained vision encoder, CLIP [20], to extract image features. These features are then used to rank all available images based on their similarity to the anchor image, resulting in a selection of the top 50 candidates that share similar semantics with the given image $I$. Next, we proceed to detect the objects in all the images, as well as the conceptual information within the questions, especially the answers. Our objective is to identify and remove images that have a high degree of overlap with the concepts mentioned in the questions and answers. To quantify this overlap, we define a semantic overlap score $s_{op}$:
@@ III-A Evaluation Setup
We pick four multi-modal large models for this evaluation.
@@
  • Otter-MPT-7B [14] shares a similar model architecture to that of Open-Flamingo but has been specifically optimized on the Multi-Modal In-Context Instruction Tuning (MIMIC-IT) dataset. This optimization aims to enhance the model’s ability to follow the in-context examples accurately for better few-shot inference.
@@
#Shots #Ans #Una BY Acc (%$\uparrow$ OoS (%$\downarrow$
@@ VI-A Visual Question Answering Datasets
VQA contributes an essential ingredient to the compelling ‘AI-Complete’ tasks. In its early emerging stage, significant efforts were dedicated to constructing a general VQA dataset, aiming to encompass all aspects of visual reasoning from textual inputs [37, 38, 39, 40]. For instance, DAQUAR [41] and COCO-QA [42] employe automated question-answer pair generation techniques, minimizing the need for extensive human labor. Visual Madlibs [43] utilizes fill-in-the-blank templates to generate VQA instances. Additionally, benchmark datasets such as [37, 44] collected large-scale human-annotated VQA data, including both natural images and abstract scenes. With increasing attention to the VQA task, the model performance on these pioneering datasets gradually approached saturation with respect to human performance.
@@ VII Conclusion and Future Work
This paper proposes a novel dataset to enable VQA models with the ability of abstention. We realize that refraining from answering questions that one does not know is a bedrock of intelligence. Therefore, testing models for this ability is important. By leveraging this dataset, we probe the robustness of multi-modal large models and identify their relatively less satisfactory performance. We then propose a straightforward approach to help train VQA models with the inclusion of unanswerable questions. We evaluate the effectiveness of this approach by integrating it into several baseline methods. The findings reported in this paper highlight two promising future research directions:
@@
2) Exploring additional possibilities for better multi-modal large model training. The existing VL large model is still in its infancy, with most efforts focused on aligning them with LLMs. However, we have discovered that these large models exhibit significant limitations when it comes to tasks such as instruction following and visual understanding. As such, a more promising direction could involve building multi-modal large models that leverage the intrinsic attributes of multi-modality, rather than relying solely on mapping it back to language.